OpenStudy (anonymous):

3(p+q)=p I've tried working this out several times, distributing the 3 and so on and cannot get one of the below answers a.q=-2/3p b.q=-3/2p b. p=-3/2q c.p=-2/3q

OpenStudy (x3_drummerchick):

are you solving for p?

OpenStudy (x3_drummerchick):

i see. 3(p+q)=p 3p + 3q = p 3q= p - 3p 3q = -2p q= - 2/3p